OpenStudy (anonymous):

Please help Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3). y = 2x + 6 y = −one half x + 3 y = −one half x + 6 y = 2x + 3

OpenStudy (mathstudent55):

Do you remember the previous problem? The first step is to find the slope of the given equation. Compare \( y = 2x + 2 \) to \(y = mx + b\) What is the slope of the given line?

OpenStudy (anonymous):

2 is the slope

OpenStudy (mathstudent55):

Great. In the previous problem we need a parallel line. Parallel lines hav the sam e slope. In this problem we need a perpendicular line. Perpendicular lines have slopes that are "negative reciprocals." All that means is that if you know the slope of a line, to find the slope of a perpendicular, write the slope of the given line as a fraction. Then flip the fraction (reciprocal) then change the sign (negative.)

OpenStudy (mathstudent55):

We know our line has a slope of 2. Let's find the slope of the perpendicular: 1. Write 2 as a fraction.

OpenStudy (anonymous):

It will be the second or third one.. right?

OpenStudy (mathstudent55):

Correct. It has to be one of those two, since the negative reciprocal of 2 is \( -\dfrac{1}{2} \).

OpenStudy (anonymous):

Is it C

OpenStudy (anonymous):

Crap I only have 2 minutes left on my test.

OpenStudy (mathstudent55):

Now we know the perpendicular line has to be: \(y = -\dfrac{1}{2}x + b\) We need b. We do the same we did in the last problem. Input the x and y of the given point into \(y = -\dfrac{1}{2}x + b\) and solve for b. \(3 = -\dfrac{1}{2}(6) + b\) \(3 = -3 + b\) \(6 = b\) Since we have \(b = 6\) we then have: \(y = -\dfrac{1}{2}x + 6\)

OpenStudy (anonymous):

YES I WAS CORRECT

OpenStudy (mathstudent55):

Great.

OpenStudy (anonymous):

Thank you so much, you were such a hug hel[

OpenStudy (anonymous):

*huge help

OpenStudy (anonymous):

ty for medal

OpenStudy (mathstudent55):

Thanks. You're welcome.